OpenStudy (hockeychick23):

Annie and her friends start a Doubles Club. The numbers below show information about all the members of the club. Boys Girls Juniors 6 8 Seniors 16 12 Express the probability that Annie chooses a girl or a junior. Use the formula: P(AUB)= P(A)+P(B)- P(AnB)

ganeshie8 (ganeshie8):

use the given formula

ganeshie8 (ganeshie8):

P(girl or junior) = P(girl) + P(junior) - P(girl and junior)

OpenStudy (hockeychick23):

Is the answer 26? thats what i got when i plugged it in

OpenStudy (hockeychick23):

I did P(AB)= P(20)+P(14)-P(8)

ganeshie8 (ganeshie8):

probability is a number between 0 and 1

ganeshie8 (ganeshie8):

P(girl) = 20/42 = ?

OpenStudy (hockeychick23):

oh ok so i have .47619+.26923- .1538461538= .59157

ganeshie8 (ganeshie8):

check P(junior) again

ganeshie8 (ganeshie8):

you should be doing : (20/42) + (14/42) - (8/42)

OpenStudy (hockeychick23):

ok now i got .6190478095